# Ali Azarbahram Shenyu Liu and Gian Paolo Incremona

### v1

@@ 4 Simulation results
While each agent maintains a Koopman matrix $K_{i}\in\mathbb{R}^{n\times n}$ which scales quadratically with $n$, this is practical in scenarios such as satellite or UAV-based sensing, where agents acquire time-varying observations under bandwidth or energy constraints. The focus in our proposed appraoch is on estimating consistent Koopman models under distributed and resource-limited conditions, rather than globally approximating nonlinear dynamics mitigating the dimensionality issue noted in Zhang and Zuazua (2023). Future extensions may incorporate scalable lifting strategies (e.g., sparse or low-rank forms) to address larger-scale problems.
Remark 6.
Algorithm 1 converges exponentially with rate governed by $\rho_{\max}$ in (15). As shown in Figure 3, approximately an accurate reconstruction is achieved after $\sim$500 iterations. In online settings, where time and resources are limited, sub-optimal iterates can still be used for short-term prediction. Figure 4 shows that even intermediate solutions yield reliable forecasts. As new data arrives, the algorithm can resume from the latest iterate, enabling continual refinement. This supports real-time deployment, where learning proceeds alongside sensing and control.
@@ 5 Conclusion
Since the matrix $\mathbf{L}^{\frac{1}{2}}$ and $I$ commute, we use Schur complement and determinantal formula Horn and Johnson (2012, Chapter 0.8.5) to conclude
@@
where we recall that the matrices $\bm{X},\tilde{\mathbf{M}}$ are defined in (9), (16). Meanwhile, it follows from the KKT condition Boyd and Vandenberghe (2004) that for any optimal solution $\mathbf{K}^{*}\in\mathbb{R}^{n\times np}$ of the constrained problem (11), there exists $\Lambda^{*}\in\mathbb{R}^{n\times np}$ such that
@@
Meanwhile, by Liu (2024, Lemma 7), $\tilde{\mathbf{M}}$ is semi-Hurwitz; that is, all its eigenvalues either have negative real parts, or are 0 and non-defective (algebraic multiplicity equals to geometric multiplicity) Additionally, because of Lemma 1 and the assumption $\alpha<\alpha_{\max}$, where $\alpha_{\max}$ is defined in (14), we either have $\lambda=0$ and non-defective so that $\lambda^{\prime}=1$ and non-defective, or $\operatorname{Re}(\lambda)<0$ and $|\lambda^{\prime}|<1$. This implies that $I+\alpha\tilde{\mathbf{M}}$ has a Jordan decomposition of the form
@@
In other words, $\mathbf{K}_{\infty},\mathbf{R}_{\infty}$ also satisfy the KKT condition of optimality for (11). Hence, $\mathbf{K}_{\infty}$ is an optimal solution for the problem (11), and $\mathbf{K}_{\infty}=\begin{bmatrix}K_{\infty}&K_{\infty}&\cdots&K_{\infty}\end{bmatrix}$, where $K_{\infty}\in\mathbb{R}^{n\times n}$ is an optimal solution for the problem (7). This completes the proof. ∎

### v2

@@ 4 Simulation results
Algorithm 1 converges exponentially with rate governed by $\rho_{\max}$ in (15). As shown in Figure 3, approximately an accurate reconstruction is achieved after $\sim$300 iterations. In online settings, where time and resources are limited, sub-optimal iterates can still be used for short-term prediction. Figure 7 shows that even intermediate solutions yield reliable forecasts. As new data arrives, the algorithm can resume from the latest iterate, enabling continual refinement. This supports real-time deployment, where learning proceeds alongside sensing and control.
Figure 8: RMSE over the prediction horizon for centralized and distributed Koopman learning methods: the proposed distributed method vs. the distributed baseline of (Hao et al., 2024a)
To contextualize the proposed method, we compare our distributed Koopman learning framework against the distributed approach introduced in (Hao et al., 2024a), which represents one of the most closely related existing results. Both approaches address the problem of learning Koopman operators from partial and distributed observations, making this comparison particularly relevant and fair. To ensure consistency across methods, the lifting functions are kept fixed and identical for the simulation runs, and both distributed algorithms are executed for the same number of iterations over the same communication graph. Moreover, since the present study focuses on autonomous prediction rather than control, the input-related components in (Hao et al., 2024a) are omitted, resulting in a formulation that aligns with our problem setup.
Prediction performance is evaluated using the root mean square error (RMSE) computed over the reserved set of unseen snapshots. Specifically, for a prediction horizon, $h=1,\ldots,9$, the RMSE is defined as
$\mathrm{RMSE}(h)=\frac{\|\hat{x}_{h}-x_{h}\|_{2}}{\sqrt{q}},$
where $\hat{x}_{h}\in\mathbb{R}^{q}$ denotes the predicted intensity field, $x_{h}$ is the corresponding ground-truth snapshot, and $q$ is the spatial dimension of the grid. This metric quantifies the average per-cell prediction error and allows for a direct comparison of multi-step forecasting accuracy across methods.
The resulting RMSE curves over the prediction horizon are shown in Figure 8. As expected, the centralized Koopman model achieves the lowest error across all horizons. Both distributed methods exhibit stable multi-step prediction behavior on unseen data, indicating that the learned operators generalize beyond the training snapshots. Among the distributed approaches, our proposed method consistently attains lower prediction error than the baseline distributed method of (Hao et al., 2024a) over the evaluated horizons, while preserving the distributed data acquisition and communication constraints.
@@ 5 Conclusion
Since the matrix $\mathbf{L}^{\frac{1}{2}}$ and $I$ commute, we use Schur complement and determinantal formula (Horn and Johnson, 2012, Chapter 0.8.5) to conclude
@@
where we recall that the matrices $\bm{X},\tilde{\mathbf{M}}$ are defined in (9), (16). Meanwhile, it follows from the KKT condition (Boyd and Vandenberghe, 2004) that for any optimal solution $\mathbf{K}^{*}\in\mathbb{R}^{n\times np}$ of the constrained problem (11a), there exists $\Lambda^{*}\in\mathbb{R}^{n\times np}$ such that
@@
Meanwhile, by (Liu, 2024, Lemma 7), $\tilde{\mathbf{M}}$ is semi-Hurwitz; that is, all its eigenvalues either have negative real parts, or are 0 and non-defective (algebraic multiplicity equals to geometric multiplicity) Additionally, because of Lemma 1 and the assumption $\alpha<\alpha_{\max}$, where $\alpha_{\max}$ is defined in (14), we either have $\lambda=0$ and non-defective so that $\lambda^{\prime}=1$ and non-defective, or $\operatorname{Re}(\lambda)<0$ and $|\lambda^{\prime}|<1$. This implies that $I+\alpha\tilde{\mathbf{M}}$ has a Jordan decomposition of the form
@@
In other words, $\mathbf{K}_{\infty},\mathbf{R}_{\infty}$ also satisfy the KKT condition of optimality for (11a). Hence, $\mathbf{K}_{\infty}$ is an optimal solution for the problem (11a), and $\mathbf{K}_{\infty}=\begin{bmatrix}K_{\infty}&K_{\infty}&\cdots&K_{\infty}\end{bmatrix}$, where $K_{\infty}\in\mathbb{R}^{n\times n}$ is an optimal solution for the problem (7). This completes the proof. ∎